OpenStudy (anonymous):

find the domain of the function g(x) = 7/6-5x

OpenStudy (bobbyleary):

Is the question 7/(6-5x) or (7/6)-5x?

OpenStudy (anonymous):

it is 7over6-5x

OpenStudy (a_clan):

All -{6/5}

OpenStudy (anonymous):

6-5x doesn't. Equal 0

OpenStudy (anonymous):

shd101wyy the answer is 0

OpenStudy (anonymous):

so if the answer is 0 this is a real number correct

OpenStudy (anonymous):

6-5x does not equal to 0

OpenStudy (bobbyleary):

To find where the function is not continuous, you set 6-5x=0 or x = (6/5). So the domain is all real numbers (any number on the number line) except for (6/5).

OpenStudy (anonymous):

bobbyleary what is the correct domain

OpenStudy (bobbyleary):

\[(-\infty,6/5)U(6/5,\infty)\] or any real number but (6/5).

OpenStudy (anonymous):

so bobby do i choose this one x is a real number x =with a line crossed through the equal sign 6/5

OpenStudy (bobbyleary):

Yes.